40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 29 22:37:54 2022, Koji, Update, IOO, WFS investigation PXL_20220730_040900843.jpgPXL_20220730_041216848.jpg
    Reply  Tue Aug 2 01:10:26 2022, Koji, Update, IOO, WFS investigation 220801_IMC_WFS_DEMOD.pdf220801_IMC_WFS_DEMOD2.pdfWFS1.pngWFS2.png
       Reply  Thu Aug 4 21:58:18 2022, Koji, Update, IOO, WFS investigation Screenshot_2022-08-04_22-10-57.pngScreenshot_2022-08-04_22-11-16.pngScreenshot_2022-08-04_22-11-53.pngScreenshot_2022-08-04_22-12-39.png
          Reply  Thu Aug 4 22:14:20 2022, Koji, Update, IOO, WFS investigation IMC_SUS_channels_TF.pdfIMC_WFS_channels_TF.pdfIMC_WFS_segment_TF.pdfIMC_WFS_220804.xlsx
             Reply  Thu Aug 4 22:14:38 2022, Koji, Update, IOO, WFS investigation Screenshot_2022-08-04_22-17-19.pngScreenshot_2022-08-04_22-18-45.png
                Reply  Thu Aug 4 22:32:31 2022, Koji, Update, IOO, Upon leaving the lab (WFS investigation) 
                   Reply  Fri Aug 5 12:42:12 2022, Koji, Update, IOO, IMC WFS: Overnight observation  Screen_Shot_2022-08-05_at_12.04.01.png
Message ID: 17053     Entry time: Tue Aug 2 01:10:26 2022     In reply to: 17048     Reply to this: 17059
Author: Koji 
Type: Update 
Category: IOO 
Subject: WFS investigation 

Continued to work on the WFS repair

Demod phase adjustment:

- Use the PDH signal to adjust the demodulation phase to have uniform signals between the segments.

- Excited laser frequency at 1234Hz by injecting 10mVpp into IMC Servo Board IN2. The input was enabled on the MC Servo screen and given the input gain of 0dB.

- Looked at the ~real time spectrum in WFS1/2 SEG1/2/3/4 I&Q after the phase rotators. Changed the demod phases 1) to have ~0deg transfer function between C1:IOO-MC_F to C1:IOO-WFSi_Ij 2) to minimize the freq signal in Q phases.
(See Attachment 1)

- Resulting change of the demod phases:

WFS1 SEG1  52.0 -> 38.0deg
WFS1 SEG2  54.0 -> 53.0deg
WFS1 SEG3  16.6 -> 33.2deg
WFS1 SEG4 103.9 ->-37.1deg

WFS2 SEG1  17.0 -> 57.8deg
WFS2 SEG2  26.6 -> 51.5deg
WFS2 SEG3  24.5 -> 44.0deg
WFS2 SEG4 -58.0 ->103.7deg

SEG4 of both WFSs had significant phase rotation. A quick check of the power spectrum indicates that the Q signals have significantly (<x1/10) lower signals (Attachment 2/3/4). So that's good.

Transfer function measurement

Now the ASCPIT/ASCYAW of the MC1/2/3 suspension were excited and the transfer functions to WFS1/2 SEG1/2/3/4 and MC Trans P/Y were measured. The analysis will come later.

Again here the Q signals have significantly lower sensitivity to the mirror motion. So it is consistent with the above observation of the spectra.

However, the quick check of the transfer functions indicated that the conventional input matrices result in the flipped dependence of the combined error signals in pitch and yaw.
This might indicate that some of the cables were not inserted into the demod board properly although the cables at the demod boards show no indication of anomaly. (See the photos in ELOG 17048)

It might be the case that the cable had been inserted with a special unusual arrangement.

In any case, this can be fixed at the input matrix. Native change of the input matrix made WFS1PIT/WFS1YAW/WFS2PIT/WFS2YAW/MC2Trans YAW servos running (after some adjustment of the servo signs).
The MC2TRANS PIT servo didn't seem to settle and run away no matter which sign is used.

It's probably better to look at the sensing matrix and figure out the proper input/output matrix carefully. So at this moment, no WFSs are working.

Note that I left the new demod phases in the system


During the transfer function measurement some filters were turned off to make the shaking smoother:

IMC ASC filters were turned off to make the FResp flat:
- MC1 ASCP/Y FM1/FM5 OFF
- MC2 ASCP/Y FM1/FM5/FM6 OFF
- MC3 ASCP/Y FM1/FM5 OFF

60Hz comb OSEM Input filters were also turned off to make the transfer functions simpler:
- MC1 INPUT FM2 OFF (60Hz comb)
- MC2 INPUT FM2 OFF (60Hz comb)
- MC3 INPUT FM2 OFF (60Hz comb)


cf. Past IMCWFS commissioning http://nodus.ligo.caltech.edu:8080/40m/12684

Attachment 1: 220801_IMC_WFS_DEMOD.pdf  63 kB  Uploaded Tue Aug 2 03:00:45 2022  | Hide | Hide all
220801_IMC_WFS_DEMOD.pdf
Attachment 2: 220801_IMC_WFS_DEMOD2.pdf  329 kB  Uploaded Tue Aug 2 03:00:49 2022  | Hide | Hide all
220801_IMC_WFS_DEMOD2.pdf
Attachment 3: WFS1.png  49 kB  Uploaded Tue Aug 2 03:46:14 2022  | Hide | Hide all
WFS1.png
Attachment 4: WFS2.png  49 kB  Uploaded Tue Aug 2 03:46:32 2022  | Hide | Hide all
WFS2.png
ELOG V3.1.3-